2x+y=5
2y=2x-8
Find the first variable and isolate it. then solve for that variable.
Solve for the second variable
Find the numerical value of the first variable
Check your solution

Answer by Alan3354(69443)

2x+y=5
2y=2x-8 ---> y = x-4
-------------
2x+y=5
2x + (x-4) = 5
3x = 9
x = 3
----
y = x-4
y = -1
=========================
This is gibberish. Which variable is the "first variable?"
---------
Find the first variable and isolate it. then solve for that variable.
Solve for the second variable
Find the numerical value of the first variable
Check your solution